OpenStudy (anonymous):

The table below shows the cube roots of different numbers: Number (x) 8 27 64 125 Cube root(y) 2 3 4 5 Part A: Does the table represent y as a function of x? Justify your answer. Part B: The total cost f(x), in dollars, for renting a bike for x hours is shown below: f(x) = 10 + 20x What is the value of f(100), and what does f(100) represent?

OpenStudy (anonymous):

@dtan5457

OpenStudy (danjs):

Y is a function of x, It tells you in the qestion y is the cubed root of x

OpenStudy (anonymous):

1've already solved part B. 1 just need help with part A

OpenStudy (danjs):

y = x^(1/3)

OpenStudy (anonymous):

ok how would you prove that though?

OpenStudy (danjs):

let y=f(x) = x^(1/3) f(8) = 8^(1/3) = 2 f(27) = 27^(1/3) = 3 . . .

OpenStudy (anonymous):

ok so the function is the proof?

OpenStudy (danjs):

You can graph the table of values, and show that for each value in the domain of y=f(x) there is one corresponding value in the range.

OpenStudy (anonymous):

Perfect thanks so much!